can you help me solve

2z -1/3 = 1/2

Answer by checkley71(8403)

2Z-1/3=1/2
2Z=1/2+1/3
2Z=(3+2)/6
2Z=5/6
Z=(5/6)/2
Z=5/6*1/2
Z=5/12 ANSWER.
PROOF:
2*5/12-1/3=1/2
10/12-1/3=1/2
(10-4)/12=1/2
6/12=1/2
1/2=1/2